OpenStudy (anonymous):

Determine the vertx of the parabola whose equation is f(x)=x^2-8x+7. Next find the minimum or maximum value. What is the x-intercepts?

OpenStudy (anonymous):

vertex=-b/2a=8/2=4 4^2-8*4+7=-9 (4,-9) the x^2 is positive so the parabola faces up so theres a minimum value minimum is the same as the vertex x intrecepts is where y=0 0=x^2-8x+7 (x-7)(x-1)=0 x=7,1 so (1,0) & (7,0)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np